How to begin in a problem like this, 4x+3/15 - 2x-3/9 = 6x+4/6 - x . Thank you for any help!
Answer by stanbon(75887)

How to begin in a problem like this,
----------------------
4x+3/15 - 2x-3/9 = 6x+4/6 - x
Multiply thru by 90 to get:
-----
6*(4x+3) - 10(2x-3) = 15(6x+4) - 90x
-----
24x + 18 -20x + 30 = 90x + 60 - 90x
-----
4x + 48 = 60
4x = 12
x = 3
================
Cheers,
Stan H.
================